OpenStudy (anonymous):

Given f(x) = the quantity of 4x plus 1, divided by 3, solve for f-1(3).

OpenStudy (anonymous):

2 4 6 8

OpenStudy (anonymous):

\[\frac{ 4x+1 }{ 3 }\] This?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

To find the Inverse, Switch the X and Y, and solve for Y.

OpenStudy (anonymous):

Your inverse should be \[\frac{ 3x-1 }{ 4 }=\] Then you plug in 3 for X in the inverse c:

OpenStudy (anonymous):

And your answer should

OpenStudy (anonymous):

be 2.

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Welcome xD